A profit of $2400 is to be divided among three people in a 7:3:2 ratio. How much should each person receive?
Answer by Deina(147)

I'm not sure about the algebraic formula for this, but it's fairly simple.
7 + 3 + 2 = 12 so
2400 / 12 = 200
7 * 200 = 1400
3 * 200 = 600
2 * 200 = 400
and bob's your uncle!